Question 11

ABC Corporation has hired you to evaluate a new FOUR year project for the firm. The project will require the purchase of a $825,400.00 work cell. Further, it will cost the firm $50,200.00 to get the work cell delivered and installed. The work cell will be straight-line depreciated to zero with a 20-year useful life. The project will require new employees to be trained at a cost of $65,300.00. The project will also use a piece of equipment the firm already owns. The equipment has been fully depreciated, but has a market value of $5,200.00. Finally, the firm will invest $10,200.00 in net working capital to ensure the project has sufficient resources to be successful.

The project will generate annual sales of $917,000.00 with expenses estimated at 37.00% of sales. Net working capital will be held constant throughout the project. The tax rate is 39.00%.

The work cell is estimated to have a market value of $464,000.00 at the end of the fourth year. The firm expects to reclaim 84.00% of the final NWC position.

The cost of capital is 15.00%.

What is the terminal cash flow for the project?
Answer #1

1)Calculation of depreciation expense and Book value at end of year 4

Depreciation expense = Cost /useful life

                       = 825400/20

                         =41270

Total accumulated depreciation over 4 years =41270 *4 = 165080

Book value of cell at end of 4 year = Cost -accumulated depreciation

                                = 825400 - 165080

                                = 660320

Gain/(loss) on sale =sale value -Book value

                = 464000 - 660320

               = - 196320 loss

Tax saving due to loss = Loss on sale *tax rate

                 = -196320 * 39%

                 =- 76564.80

After tax sale value= sale value - Tax saving

                      = 464000 - (-76564.80)

                     = 464000 + 76564.80

                       = 540564.80

Calculation of terminal value :

After tax sale value 540564.80
Net working capital released (10200*84%) 8568
Terminal value of project 549132.80